# MARS: A Multi-Agent Framework Incorporating Socratic Guidance for Automated Prompt Optimization

### v1

@@ Appendix G Universal Optimum Solution
Manager
You are the ChatManager, please schedule the right Agent for the job as needed.
UserProxy
Here is a topic for geometric graph generation: Given a full SVG path element containing multiple commands, determine the geometric shape that would be generated if one were to execute the full path element. For example: This SVG path element ‘<path d="M 64.00,63.00 L 44.00,63.00 L 44.00,50.00 L 64.00,50.00 L 64.00,45.00 L 85.00,57.00 L 64.00,68.00 L 64.00,63.00"/>’ draws a: Options: (A) circle (B) heptagon (C) hexagon (D) kite (E) line (F) octagon (G) pentagon (H) rectangle (I) sector (J) triangle I want to input a prompt and this topic into the big language model so that the big language model outputs the highest correctness rate. Please generate the most suitable prompt according to the requirements I just mentioned.
@@
Split the task ’Here is a topic for geometric graph generation: Given a full SVG path element containing multiple commands, determine the geometric shape that would be generated if one were to execute the full path element. For example: This SVG path element <path d=M 64.00,63.00 L 44.00,63.00 L 44.00,50.00 L 64.00,50.00 L 64.00,45.00 L 85.00,57.00 L 64.00,68.00 L 64.00,63.00""/> draws a Options: (A) circle (B) heptagon (C) hexagon (D) kite (E) line (F) octagon (G) pentagon (H) rectangle (I) sector (J) triangle I want to input a prompt and this topic into the big language model so that the big language model outputs the highest correctness rate. Please generate the most suitable prompt according to the requirements I just mentioned.’ into detailed steps and details. For example, for the clinical medicine Test, the task is planned as follows: Total steps: 4 Step 1: Analyze the input requirements, focusing on the type of clinical medicine question and the format of the options. Step 2: Design a prompt that encourages the model to consider the specific clinical characteristics of the condition described in the question and match the most appropriate option based on medical knowledge. Step 3: Request the model to evaluate each option in the context of clinical presentation, symptoms, and diagnostic characteristics of the condition to ensure it selects the most accurate answer. Step 4: Test and refine the prompt to ensure the model produces the highest correctness rate for similar clinical medicine questions.
@@
You are an evaluator responsible for judging the correctness of a given task. Your output must strictly follow these rules: 1. If the task is judged as correct, output only: [True] 2. If the task is judged as incorrect, output: [False] [suggestion: <reason for the incorrect judgment>] Replace ‘<reason for the incorrect judgment>‘ with a clear and concise explanation of why the task is incorrect. Do not include any additional text, comments, or explanations beyond the specified format.
@@
Prompt: Systematically analyze the given SVG path element by first breaking it down into its individual commands, such as ’M’ (move to), ’L’ (line to), and others. For each command, map the sequence of points it generates, ensuring you accurately trace the path step by step. As you follow the path, focus on identifying key geometric properties, such as equal side lengths, parallel lines, specific angles, or symmetries, that emerge between consecutive points. Use these properties to classify the shape based on its defining characteristics. For example, given the path <path d=M 64.00,63.00 L 44.00,63.00 L 44.00,50.00 L 64.00,50.00 L 64.00,45.00 L 85.00,57.00 L 64.00,68.00 L 64.00,63.00/>, calculate the distances between points to check for equal side lengths, measure angles to identify parallelism or perpendicularity, and look for symmetries that align with known geometric shapes. Based on these observations, determine whether the shape is a circle, heptagon, hexagon, kite, line, octagon, pentagon, rectangle, sector, or triangle. Provide the correct answer by reasoning through the geometric properties derived from the path commands. Question: question Please don’t output the process of doing the question, only the content of the answer.The answer should be a parenthesis containing the capital letter of the chosen answer. please do not add any other spaces or symbols.
@@
Think step by step to solve linguistically diverse elementary school math application problems. Break down the problem into 2-8 logical steps, perform the necessary calculations at each step, and provide the final result. Ensure accuracy by carefully following the problem’s instructions and verifying each intermediate step. For example: Input: Janet’s ducks lay 16 eggs per day. She eats three for breakfast every morning and bakes muffins for her friends every day with four. She sells the remainder at the farmers’ market daily for \$2 per fresh duck egg. How much in dollars does she make every day at the farmers’ market? Step 1: Calculate the total eggs used daily: 3 (eaten) + 4 (baked) = 7 eggs. Step 2: Subtract the used eggs from the total laid: 16 - 7 = 9 eggs. Step 3: Calculate the daily earnings: 9 eggs × $2=$18. Answer: 18 Follow this structured approach to solve similar problems.

### v2

@@ Acknowledgments
Split the task ’Here is a topic for geometric graph generation: Given a full SVG path element containing multiple commands, determine the geometric shape that would be generated if one were to execute the full path element. For example: This SVG path element ¡path d=M 64.00,63.00 L 44.00,63.00 L 44.00,50.00 L 64.00,50.00 L 64.00,45.00 L 85.00,57.00 L 64.00,68.00 L 64.00,63.00””/¿ draws a Options: (A) circle (B) heptagon (C) hexagon (D) kite (E) line (F) octagon (G) pentagon (H) rectangle (I) sector (J) triangle I want to input a prompt and this topic into the big language model so that the big language model outputs the highest correctness rate. Please generate the most suitable prompt according to the requirements I just mentioned.’ into detailed steps and details. For example, for the clinical medicine Test, the task is planned as follows: Total steps: 4 Step 1: Analyze the input requirements, focusing on the type of clinical medicine question and the format of the options. Step 2: Design a prompt that encourages the model to consider the specific clinical characteristics of the condition described in the question and match the most appropriate option based on medical knowledge. Step 3: Request the model to evaluate each option in the context of clinical presentation, symptoms, and diagnostic characteristics of the condition to ensure it selects the most accurate answer. Step 4: Test and refine the prompt to ensure the model produces the highest correctness rate for similar clinical medicine questions.
@@
You are an evaluator responsible for judging the correctness of a given task. Your output must strictly follow these rules: 1. If the task is judged as correct, output only: [True] 2. If the task is judged as incorrect, output: [False] [suggestion: ¡reason for the incorrect judgment¿] Replace ‘¡reason for the incorrect judgment¿‘ with a clear and concise explanation of why the task is incorrect. Do not include any additional text, comments, or explanations beyond the specified format.
@@
Prompt: Systematically analyze the given SVG path element by first breaking it down into its individual commands, such as ’M’ (move to), ’L’ (line to), and others. For each command, map the sequence of points it generates, ensuring you accurately trace the path step by step. As you follow the path, focus on identifying key geometric properties, such as equal side lengths, parallel lines, specific angles, or symmetries, that emerge between consecutive points. Use these properties to classify the shape based on its defining characteristics. For example, given the path ¡path d=M 64.00,63.00 L 44.00,63.00 L 44.00,50.00 L 64.00,50.00 L 64.00,45.00 L 85.00,57.00 L 64.00,68.00 L 64.00,63.00/¿, calculate the distances between points to check for equal side lengths, measure angles to identify parallelism or perpendicularity, and look for symmetries that align with known geometric shapes. Based on these observations, determine whether the shape is a circle, heptagon, hexagon, kite, line, octagon, pentagon, rectangle, sector, or triangle. Provide the correct answer by reasoning through the geometric properties derived from the path commands. Question: question Please don’t output the process of doing the question, only the content of the answer.The answer should be a parenthesis containing the capital letter of the chosen answer. please do not add any other spaces or symbols.
@@
Think step by step to solve linguistically diverse elementary school math application problems. Break down the problem into 2-8 logical steps, perform the necessary calculations at each step, and provide the final result. Ensure accuracy by carefully following the problem’s instructions and verifying each intermediate step. For example: Input: Janet’s ducks lay 16 eggs per day. She eats three for breakfast every morning and bakes muffins for her friends every day with four. She sells the remainder at the farmers’ market daily for \$2 per fresh duck egg. How much in dollars does she make every day at the farmers’ market? Step 1: Calculate the total eggs used daily: 3 (eaten) + 4 (baked) = 7 eggs. Step 2: Subtract the used eggs from the total laid: 16 - 7 = 9 eggs. Step 3: Calculate the daily earnings: 9 eggs × 2=2=18. Answer: 18 Follow this structured approach to solve similar problems.